# Rheological Analysis and Evaluation of Measurement Techniques for the Curing Polymethylmethacrylate Bone Cement in Vertebroplasty

### v1

@@ 3.7 Validity of the Cox-Merz rule for PMMA bone cement
According to the Cox-Merz rule, for the same magnitude of shear rate and angular frequency, the respectively measured steady shear viscosity and complex viscosity must also be equal 26. To extend this rule to materials with yield stress, Doraiswamy et al. 39 suggested multiplying the angular frequency to the strain amplitude for the Cox-Merz rule to be true. To validate this for the Vertecem V+ bone cement, the results from shear rate sweep and three repetitions of frequency sweep measurements at various strain amplitudes are superimposed in Figure 3(a). Note that, as we learned from the amplitude sweep measurements shown in Figure 2(c), the 0.01% strain amplitude is within the linear viscoelastic region, 0.03% is near the transition to non-linear region, 0.2% is in the non-linear region. The frequency sweep curves for 0.01% did not coincide with those of steady shear viscosity, but they were closer for 0.03%. The graphs coincided at 0.2%, which is in the non-linear viscoelastic range. As a check, a single frequency sweep was carried out at 200% strain amplitude, which is beyond the crossover point. The resulting curve seemed to extrapolate the curves from the tests with 0.2% strain, and coincided with the rotational shear rate graph until the aforementioned dip in viscosity due to wall slip. This is surprising since the frequency sweep is generally done at a strain amplitude that lies within the linear viscoelastic range. A possible explanation is provided by Shafiei-Sabet et al. 40, that the PMMA structures formed as a result of polymerization break under shear flow, but oscillatory tests at low strains do not affect these structures. This means that the Cox-Merz rule is not applicable to the Vertecem bone cement if the strain amplitude is low, i.e. within the linear viscoelastic range, but it is only valid for higher strain amplitudes in the non-linear range. This reveals the conditional nature of the validity of the Cox-Merz rule for the Vertecem bone cement.
@@ 3.8 Optical microscopy and visual inspection
Examples of cured bone cement samples after tests on the rheometer are shown in Figure 6. These samples were inspected microscopically to gain better insight into the results. Figure 6a shows an unsheared sample simply left at rest between the rheometer plates until it was cured, from test Rh9. Figure 6b shows sample subjected to 100 s for 15 minutes before letting it cure, from test Rh10. We observed circular “ridge”-like formations formed in the sheared sample as a result of continuous rotational deformation. The images from optical microscopy, shown in Figure 6e and 6f, showed that the depth of the ridges went almost through the entire thickness of the sample. The same ridges were observed even in the sample sheared at 0.1 s for 20 minutes (referred to test Rh8, Figure 3(c)), as shown in Figure 6c. A closer look, shown in Figure 6g, revealed that the sample looked normal in the center upto a radius of about 13.66 mm, and the ridges were formed only beyond this radius. In a parallel plate rheometer, the shear rate is not uniform across the sample, hence the specified shear rate is actually the average shear rate along the radius. Given that the shear rate at the periphery is 1.5 times the specified shear rate, and the shear rate increases linearly with the radius, we could calculate the critical shear rate as 0.082 s, beyond which the ridge formation was observed.
Interestingly, there were no visible ridges in the sample exposed to oscillatory deformation with 20% strain and 1 Hz for 30 minutes (refer test Rh1b, Figure 1(b)), as shown in Figure 6d. This implies that rotational shearing on the rheometer causes deformations in the bone cement that results in the formation of deep circular ridges, likely stemming from the combination of elastic-like instabilities and possibly Taylor-type vortex flows, as has been observed for other types of materials 41, 42, 43. This could happen even at low shear rates if the sample is sheared for sufficiently long time. The ridges would not only cause a loss of contact with the top plate, but also a lack of sufficient contact between neighbouring flow layers of the bone cement needed for viscous resistance, which would result in underestimation of the viscosity. This phenomenon limits the conditions in which the rheometer tests can be done on the bone cement and thus makes standardization of tests very difficult.
@@ 3.9 Tests on the custom-made injector (Tests Inj1 – Inj5)
The results of the injection experiments using the custom-made injector setup at various flow rates are presented in Figure 7, which shows the measured force against time normalized by the total duration of injection. As expected, the results showed that the higher the flow rate, the more injection force is required. For the two slowest flow rates, the force gradually increased because the duration of injection was long enough and the shear rates were small enough for the cement to cure while being injected. To obtain reliable results, the average force values for each flow rate over a 5 second time period closest to the 245 seconds mark from the time of mixing were used for the regression analysis, as shown in Figure 6(b). This was done to avoid any influence of curing time on the outcome of curve-fitting. The regression was carried out using Equations 5 and 6 obtained from analytical calculations, with power law parameters and as fitting parameters. The regression analysis gave a good fit with yielding values Pa s and . The value of is similar to those observed in test Rh7a for shear rate less than 1 s and in test Rh8 during the initial 30 seconds. Using Equation 4, we could calculate the maximum shear rate, which were 10 s and 165 s for the lowest and the highest flow rates respectively. For shear rates in this range, we observed decreasing viscosity with time in test Rh8 (Figure 3(c)). However, if this was truly the case, it would cause a decrease in the injection force with time, which was not observed in our results further evidence that the reducing viscosity seen in Figure 3(c) is due to the ridges formed in the sample as a result of the rotational deformation.
@@ 3.10 Comparison of the methods
| Test | Angular frequency * Strain amplitude/ | (Pa s) | |
@@
The values of the parameters of the power law equation and are useful for analyzing and simulating the injection flow behaviour of bone cement. Usually, these values are extracted from experimentally measured flow curves using rheometer tests like test Rh7a, or the oscillatory frequency sweep test Rh6 if the Cox-Merz rule is known to be valid. They can also be extracted from a capillary rheometer or an injector setup like in tests Inj1–5. The power law parameters extracted from these tests are summarized in Table 2.
@@
Analytical calculations show that the shear rate range during an actual injection could go as high as 200 s. As previously discussed, the rotational rheometer is not suitable at such high shear rates because of the wall sip and ridge formation caused by rotational deformation. This could be avoided by reducing the gap size between the rheometer plates, but that cannot be done in this case since that the gap size must also be sufficiently large to ensure no particle effects. For the same reason, we could not use our cone-plate setup on the rheometer as it could only be used with a fixed gap of 49 microns, which is not large enough for our application. Hence, only measurements for shear rates below 1 s can be reliably used, as the power law parameters thus obtained are close to those obtained from the injection experiment. It must be noted, however, that even for low shear rates, the ridge formation could occur if the test duration was sufficiently long. It is important to exercise these These cautions when performing rotational rheometer tests on PMMA bone cements.
@@
The flow curves can also be obtained from the oscillatory frequency sweep, but as seen from Figure 5, it is important to ensure that the strain amplitude used for the frequency sweep lies within the non-linear viscoelastic range. Only then is the Cox-Merz rule valid and a feasible comparison to steady shear flow (like during an injection) possible. This is a crucial finding since this is in contrast to the conventional norm of performing the frequency sweep with a strain amplitude within the linear viscoelastic region. Note that the viscosity drop seen as a result of wall slip in the steady shear viscosity from rotational measurements is not seen in the oscillatory frequency sweep results. This, along with previously detailed visual and microscopical inspection, confirmed that that problem of wall slip or ridge formation is absent in oscillatory measurements. This demonstrates the advantage of oscillatory tests over rotational tests, especially as the shear rates go higher.
@@
The setup of a capillary rheometer or the custom-made injector setup used in this study is closest to an actual vertebroplasty setting. This kind of setup circumvents the drawbacks of the rotational or oscillatory rheometry, and yields mechanical parameters like effective or apparent viscosity and injection force that is most representative of the real-world setting. It is important to note that the flow model used and the analytical calculations play an important role in the accuracy of the extracted rheological parameters. The Herschel-Buckley model is arguably the more suited model since the bone cement exhibited yield stress. However, the yield stress is low compared to the shear stresses applied during the measurements in most cases, hence the power law model was used in this study for simplicity. There needs to be correction factors applied for flow entry, exit, change in tube diameter, etc., but they are beyond the scope of this work. Hence, the rheological parameters extracted from our setup are indicative numbers, and are not expected to have the accuracy or sophistication of industry-standard instruments.
@@ 4 Conclusion
From the various tests done, the curing PMMA bone cement was found to exhibit viscoelastic behaviour with a two-phase curing process: a slow initial phase followed by a rapid curing phase. The curing process is significantly affected by deformations and surrounding temperature. Exposure to physiological body temperature immediately triggered the rapid curing phase resulting in sudden increase in viscosity. This could negatively affect the vertebroplasty procedure if it is done in step-wise injections as a result of flow instabilities caused by injection of a low-viscosity fluid in a high-viscosity fluid environment. Properties like low yield stress and fluid-like behaviour under deformation, as opposed to predominantly elastic contribution at rest, make it suitable for injection.
While measuring the shear-thinning characteristics, rotational rheometry underestimated the viscosity above 0.082 shear rate. It was confirmed from microscopical visual inspection that the dip in viscosity was caused due to slippage between material and rheometer plates, and formation of circular ridges within the sample material. The same was not observed in a sample subjected to oscillatory rheometry and tests using our injector setup.
An important finding of the study is that Cox-Merz rule is valid only for oscillatory frequency sweep tests done at strain amplitude in the non-linear viscoelastic region, in contrast to the conventional norm of doing the test within the linear viscoelastic region. This conditionally validity of the Cox-Merz rule is important to take into consideration, especially when oscillatory measurements are used to obtain the flow curves.

### v2

@@ 3.7 Validity of the Cox-Merz rule for PMMA bone cement
According to the Cox-Merz rule, for the same magnitude of shear rate and angular frequency, the respectively measured steady shear viscosity and complex viscosity must also be equal 37. To extend this rule to materials with yield stress, Doraiswamy et al. 47 suggested multiplying the angular frequency by the strain amplitude for the Cox-Merz rule to be true. To validate this for the Vertecem V+ bone cement, the results from the shear rate sweep and three repetitions of frequency sweep measurements at various strain amplitudes are superimposed in Figure 5. Note that, as we learned from the amplitude sweep measurements shown in Figure 3(c), the 0.01% strain amplitude is within the linear viscoelastic region, 0.03% is near the transition to non-linear region, 0.2% is in the non-linear region. The frequency sweep curves for 0.01% did not coincide with those of steady shear viscosity, but the same were closer for 0.03%. The graphs coincided at 0.2%, i.e. when the applied strain amplitude was in the non-linear viscoelastic range. As a check, a single frequency sweep was carried out at 200% strain amplitude, which is beyond the crossover point. The resulting curve seemed to extrapolate the curves from the tests with 0.2% strain, and coincided with the rotational shear rate graph until the aforementioned dip in viscosity due to wall slip. This is surprising, since conventionally, the frequency sweep is carried out at a strain amplitude lying within the linear viscoelastic range. However, our measurements show that the Cox-Merz rule is indeed valid, but only in the non-linear viscoelastic regime. A possible explanation is provided by Shafiei-Sabet et al. 48, which says that the PMMA structures formed as a result of polymerization break under rotational shear flow, but the deformations during the oscillatory tests within the linear viscoelastic region are too low to affect these structures. Hence, the curves of the rotational tests and the oscillatory tests do not coincide in the linear viscoelastic region, but as the strain amplitude goes higher in the non-linear regime, the polymer structures break like in the case of rotational shear flow, and the curves coincide. This reveals that the validity of the Cox-Merz rule for the Vertecem bone cement is conditional, i.e. only in the non-linear viscoelastic regime.
@@ 3.8 Optical microscopy and visual inspection
Examples of cured bone cement samples after tests on the rheometer are shown in Figure 6. These samples were inspected microscopically to gain better insight into the results. Figure 6a shows an unsheared sample simply left at rest between the rheometer plates until it was cured, from test Rh9. Figure 6b shows a sample subjected to 100 s-1 for 15 minutes before letting it cure, from test Rh10. We observed circular “ridge”-like formations formed in the sheared sample as a result of continuous rotational deformation. The images from optical microscopy, shown in Figure 6e and 6f, showed that the depth of the ridges went almost through the entire thickness of the sample. The same ridges were observed even in the sample sheared at 0.1 s-1 for 20 minutes (referred to test Rh8, Figure 4(c)), as shown in Figure 6c. A closer look, shown in Figure 6g, revealed that the sample looked normal in the center up to a radius of about 13.66 mm, and the ridges were formed only beyond this radius. In a parallel plate rheometer, the shear rate is not uniform across the sample, hence the specified shear rate is actually the average shear rate along the radius. Given that the shear rate at the periphery is 1.5 times the specified shear rate, and the shear rate increases linearly with the radius, we could calculate the critical shear rate as 0.082 s-1, beyond which the ridge formation was observed.
Interestingly, there were no visible ridges in the sample exposed to oscillatory deformation with 20% strain and 1 Hz for 30 minutes (refer test Rh1b, Figure 2(b)), as shown in Figure 6d. No net deformation and slip occurs because of the reversible nature of the oscillatory torsion. On the other hand, rotational shearing on the rheometer causes deformations in the bone cement that result in the formation of deep circular ridges, likely stemming from the combination of elastic-like instabilities and possibly Taylor-type vortex flows, as has been observed for other types of materials 49, 50, 51. This could happen even at low shear rates if the sample is sheared for a sufficiently long time. The ridges would not only cause a loss of contact with the top plate, but also a lack of sufficient contact between adjacent flow layers of the bone cement needed for viscous resistance, which would result in underestimation of the viscosity, as has been observed in our rotational rheometer measurements. Hence, this phenomenon limits the conditions in which the rotational rheometry can be done on the bone cement.
@@ 3.9 Tests on the custom-made injector (Tests Inj1 – Inj5)
The results of the injection experiments using the custom-made injector setup at various flow rates are presented in Figure 7, wherein Figure 7(a) shows the measured force against time normalized by the total duration of injection. As expected, the results showed that the higher the flow rate, the more injection force is required. For the two slowest flow rates, the force gradually increased. This was probably because the duration of injection was long enough and the shear rates were small enough for the cement to cure while being injected. To obtain reliable results, the average force values for each flow rate over a 5-second time period closest to the 245-second mark from the time of mixing were used for the regression analysis, as shown in Figure 7(b). This was done to avoid any influence of curing time on the outcome of curve-fitting. The regression was carried out using Equations 5 and 6 obtained from analytical calculations, with power law parameters and as fitting parameters. The regression analysis gave a good fit with yielding values Pa sn and . The value of is similar to those observed in test Rh7a for shear rate less than 1 s-1 and in test Rh8 during the initial 30 seconds. We confirmed no difference between the fitting of power law and Herschley-Bulkey models in this case (refer Figure S4 in supporting information).
Using Equation 4, we could calculate the maximum shear rates, which were 10 s-1 and 165 s-1 for the lowest and the highest flow rates respectively. For shear rates in this range, we observed decreasing viscosity with time in test Rh8 (Figure 4(c)). However, if this was truly the case, it would cause a decrease in the injection force with time, which was not observed in our results. This is further evidence that the reducing viscosity seen in Figure 4(c) is due to the ridges formed in the sample as a result of the rotational deformation. In fact, surface degradation due to wall slip in injection flow conditions has been observed in previous studies 52, 53. However, there was no such evidence of wall slip in our experiments, as no difference in the surface consistencies was observed in the bone cement samples injected at different flow rates, despite the much higher shear rates compared to that of 0.082 s-1 observed in the rotational rheometry. This is probably because the injection flow setting allows for much higher friction between the flowing bone cement and the walls as a result of the high injection pressures. The same was not possible in rotational rheometry with parallel plates, as even very small normal force caused the bone cement to squeeze out from between the parallel plates during our trials with normal force control.
@@ 3.10 Comparison of the methods
| Test | Angular frequency * Strain amplitude/ | (Pa sn) | |
@@
The values of the parameters of the power law equation and are useful for analyzing and simulating the injection flow behavior of bone cement. Usually, these values are extracted from experimentally measured flow curves using rheometer tests like test Rh7a, or the oscillatory frequency sweep test Rh6 if the Cox-Merz rule is known to be valid. They can also be extracted from a capillary rheometer or an injector setup like in tests Inj1–5. The power law parameters extracted from these tests are summarized in Table 2.
@@
Analytical calculations show that the shear rate range during an actual injection could go as high as 200 s-1. As previously discussed, the rotational rheometer is not suitable at such high shear rates because of the wall sip and ridge formation caused by rotational deformation. This could be avoided by reducing the gap size between the rheometer plates, but that cannot be done in this case since the gap size must also be sufficiently large to ensure no particle effects. For the same reason, we could not use our cone-plate setup on the rheometer as it could only be used with a fixed gap of 49 microns, which is not large enough for our application. Hence, only measurements for shear rates below 1 s-1 can be reliably used, as the power law parameters thus obtained are close to those obtained from the injection experiment. It must be noted, however, that even for low shear rates, the ridge formation could occur if the test duration was sufficiently long. It is important to exercise these cautions when performing rotational rheometer tests on PMMA bone cements.
@@
The flow curves can also be obtained from the oscillatory frequency sweep, but as seen from Figure 5, it is important to ensure that the strain amplitude used for the frequency sweep lies within the non-linear viscoelastic region. Only then is the Cox-Merz rule valid and a feasible comparison to steady shear flow possible. This is a crucial finding, since this is in contrast to the conventional norm of performing the frequency sweep with a strain amplitude within the linear viscoelastic region. Note that the viscosity drop seen as a result of wall slip in the steady shear viscosity from rotational measurements is not seen in the oscillatory frequency sweep results. This, along with previously detailed visual and microscopical inspection, confirmed that the problem of wall slip or ridge formation is absent in oscillatory measurements. This demonstrates the advantage of oscillatory tests over rotational tests, especially as the shear rates go higher.
@@
The setup of a capillary rheometer or the custom-made injector setup used in this study is closest to an actual vertebroplasty setting. This kind of setup circumvents the drawbacks of the rotational or oscillatory rheometry, and yields mechanical parameters like effective or apparent viscosity and injection force that is most representative of the real-world conditions. It is important to note that the flow model used and the analytical calculations play an important role in the accuracy of the extracted rheological parameters. The Herschel-Buckley model is generally used for materials with yield stress. However, we did not observe significant difference in the fitting compared to the power law model (refer Figures S2, S3, and S4 in supporting information). Therefore, the power law model was used in this study for simplicity. As a drawback of this study, for an experimental setup involving injection through a circular tube or capillary, there need to be correction factors applied for flow entry, exit, change in tube diameter, etc. However, they have been ignored here since they are beyond the scope of this work. Hence, the rheological parameters extracted from our injector setup should be regarded as indicative and dependent on the specific setup and experimental conditions.
@@ 4 Conclusion
In this study, we conducted various tests using rotational rheometry, oscillatory rheometry, and a custom-made injector setup to understand the rheological behavior of a PMMA bone cement, here Vertecem V+, in the context of vertebroplasty. The bone cement exhibited low yield stress and fluid-like behavior under deformation, despite predominantly elastic contribution at rest, making it suitable for injection. Shear-thinning characteristics were observed, with the power law model parameters ranging between Pa sn and across measurements. The two-phase curing process allowed the bone cement to be worked and injected during the initial slow curing phase, followed by the rapid curing phase. The curing rate of the bone cement was diminished upon deformation, while it rapidly accelerated upon exposure to the human body temperature. This needs to be considered when injecting bone cement in steps of smaller injections during vertebroplasty. Rotational rheometry was found to be susceptible to measurement artifacts like underestimation of viscosity, especially above 0.082 shear rate. The cause was confirmed to be slippage between the bone cement sample and rheometer plates, and formation of circular ridges within the sample material. These measurement artifacts were not observed in oscillatory rheometry and injection tests. Finally, the Cox-Merz rule was found to be valid only for strain amplitudes in the non-linear viscoelastic region. This conditional validity of the Cox-Merz rule is critical when using oscillatory measurements to obtain the shear-thinning characteristics. Our findings underscore the impact of conditions and measurement methods on measured flow behavior. Understanding these nuances is crucial when interpreting bone cement rheology data for medical applications or their simulations.
@@
CRediT statement
Zubin Trivedi: Conceptualization, Methodology, Formal analysis, Investigation, Data curation, Validation, Writing - Original Draft, Writing – review & editing, Visualization Jacek K. Wychowaniec: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Validation, Visualization, Writing – original draft, Writing – review & editing Dominic Gehweiler: Methodology, Software, Resources, Project administration Christoph M. Sprecher: Methodology, Investigation Andreas Boger: Formal analysis, Writing - Review & Editing, Supervision Boyko Gueorguiev: Project administration, Supervision, Funding acquisition Matteo D’Este: Resources, Writing - Review & Editing, Supervision, Project administration Tim Ricken: Supervision, Project administration, Funding acquisition Oliver Röhrle: Supervision, Project administration, Funding acquisition
Additional figures containing the following supporting information are provided.
Results of benchmark test at 3 mNm maximum torque, 1 Hz frequency, 23 °C on bone cement prepared by our mixing process
Comparison of fitting using the Herschel-Bulkley and the power law rheological models on the rotational shear rate sweep test from 10-4 to 10-2 s-1
Comparison of fitting using the Herschel-Bulkley and the power law rheological models on the rotational shear rate sweep test from 10-2 to 1 s-1
Comparison of fitting using the Herschel-Bulkley and the power law rheological models on the injection force measurement tests at various flow rates
Measurement data and test settings can be accessed from our data repository through the link https://doi.org/10.18419/darus-4004.